OpenStudy (anonymous):

What is the simplified form of x minus 4 over x squared minus x minus 12 • x minus 4 over x squared minus 8x plus 16 ?

OpenStudy (anonymous):

\[\frac{x-4 }{ x^2}-12*x- \frac{ 4 }{ x^2 }-8x+16\]

OpenStudy (anonymous):

There is it set up, now, what can you do first to simplify?

OpenStudy (anonymous):

Also congrats on your first question in mathematics.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

so, first, I should multiply12 and x, so I get 12x, and then I combine like terms, right? so, simplified it would be \[(x-8)-20x+16\]right?

OpenStudy (anonymous):

You cannot combine 12 and x.

OpenStudy (anonymous):

I can't? I thought it that when you multiply a number and an unknown, you get \[number(unknown)\]

OpenStudy (anonymous):

\[\frac{ x-8 }{ x^2 }+2x+16\]